# EarthGPT-X: Enabling MLLMs to Flexibly and Comprehensively Understand Multi-Source Remote Sensing Imagery

### v1

@@ VI Conclusion
In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources. As the first RS spatial model to unify referring and grounding tasks within a visual prompting framework, EarthGPT-X represents a significant advancement in the application of MLLMs to the RS domain. Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery. Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios, such as SAR and infrared. Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-grained visual understanding at the image, region, pixel levels, etc., allowing for precise and intelligent analyses in real-world applications. In the future, we plan to enrich the SAR and infrared datasets to encompass a broader range of tasks, addressing the current limitation of narrow scope. Additionally, we aim to enable EarthGPT-X to perform more complex visual reasoning tasks across multiple visual modalities, further solidifying EarthGPT-X as a versatile and powerful tool for comprehensive RS imagery analyses.

### v2

@@ VII Conclusion
In this paper, we propose EarthGPT-X, a flexible and versatile MLLM specifically designed for understanding RS imagery from multiple sources. Technically, the hybrid signal mutual understanding mechanism is developed to enhance the interplay between dense visual features, sparse free-form visual prompts, and text instructions, facilitating a multi-grained understanding of complex RS imagery. Furthermore, a novel cross-domain one-stage fusion training strategy is proposed to generalize general-domain knowledge into RS multi-source scenarios. Through the designed visual prompt learning framework and training strategy, EarthGPT-X achieves multi-level visual understanding at the image, region, and pixel levels, etc., allowing for precise and intelligent analyses in real-world applications. In the future, we aim to expand EarthGPT-X towards a wider range of tasks and more advanced visual reasoning capabilities across diverse modalities, ultimately strengthening its versatility and effectiveness in comprehensive RS imagery analysis.